OpenStudy (anonymous):

what happens to the force of attraction between two objects if the masses of both the object is doubled?

OpenStudy (anonymous):

I suppose you mean Gravity Force :) F=Gm1m2/r^2 If we double the mass: F'=G*2m1*2m2/r^2 We can cancel G and r^2 F/F'=m1m2/4m1m2 F/F'=1/4 F'=4F The force will be 4 times larger :)

OpenStudy (anonymous):

could you please tell me the distance between those two objects,i can not simply regard it as a particle without the distance and their volume

OpenStudy (anonymous):

i did nt understand wat u asked 4 @shunyao ....but i think @imron07 's answer is right

OpenStudy (anonymous):

you know you can not just simply regard those two objects as two particles and use Newton's equation

OpenStudy (anonymous):

particle????

OpenStudy (anonymous):

yes,you don't know that?

OpenStudy (anonymous):

nop.......sry

OpenStudy (anonymous):

the Newton's formula can not be used in every situation,for objects that can not be regarded as particles,you must use Calculus to solve it

OpenStudy (anonymous):

Yes, @shunyao is somewhat right. Even if Newton's Law of Gravity hold for every object, the answer of this question won't be that simple if the objects dimension are not simple. It only hold if: The two object are particle. The two object are sphere with 'nicely' distributed density. Like planets and sun

OpenStudy (anonymous):

oooo.......